OpenStudy (anonymous):

If A=2/3 B, B=2/3 C, and C=2/3 D, what part of D is in B? a-8/27 b-4/9 c-2/3 d-75 e-4/3 Do I just multiply 2/3 by 2/3?

OpenStudy (anonymous):

You could solve for a variable and insert it into the equation. The question, to me, appears to require simple combinations. First, I suppose you would let B=2/3*2/3*D (substitute C for 2/3*d) Second, solve for D. You should get 9/4B. Third, You need to remove the B. Set be equal to 2/3C. That should give you 8/27. I hope that is right. I just jotted that down while working on a project. Good luck.

OpenStudy (anonymous):

Oh, thank you..

OpenStudy (anonymous):

Hold on. I just saw that you wrote A=2/3b

OpenStudy (anonymous):

Ok. Here it is D=3/2*C then D=3/2*3/2*B (because C=3/2B). Following that, D=3/2*3/2*3/2*A. That means that A=8/27*D and B=3/2*A. Plug in for a and you get that B=4/9D. The question is a bit confusing, but it would be 4/9 parts of D is equal or into B. You can also use a Venn Diagram to find the intersection of B and D this leads to the appropriate answer.